why at the same temperature and pressure all gases will possess same KE per mol?

Grade:11

1 Answers

Mahendarsingh Mahendarsingh
6 years ago
let the initial pressure and volume of two vessel be p and v. and the final pressure will be p’the number of moles in the system will remain constantn1+n2=n1’+n2’pv/RT1 +pv/RT2 =p’v/RT +p’v/RTfrom this p’ will bep’=p(1/T1+1/T2)*T/2 -------------(1)now as the change in volume is zero and assuming the loss in surrounding is zero the internal energy will be constantinternal energy before= internal energy after joiningCvPV/R +CvPV/R=CvP’(V1+V2)/RP’=P--------(2)Using 1 and 2T=2T1T2/(T1+T2)